OpenStudy (anonymous):

Verify each identity: (secx-tanx)^2=(1-sinx)/(1+sinx)

OpenStudy (ranga):

(secx-tanx)^2 = { 1/cos(x) - sin(x)/cos(x) }^2 = { (1-sin(x)) }^2 / cos^2(x) = { (1-sin(x)) }^2 / { 1 - sin^2(x) } = { (1-sin(x)) }^2 / { (1 + sin(x)) * (1 -sin(x)) } = {1-sin(x)}/{1+sin(x)}

OpenStudy (anonymous):

ok :) i understand most of it except for the last part when you multiplied the bottom by (1-sinx)

OpenStudy (ranga):

a^2 - b^2 = (a+b) * (a-b) 1 - sin^2(x) = 1^2 - sin^2(x) = (1+sin(x)) * (1-sin(x))

OpenStudy (anonymous):

how come you didn't multiply the top?

OpenStudy (ranga):

We are not multiplying the top and bottom with the same thing here. We are just factoring the denominator, { 1 - sin^2(x) }, as (1+sin(x)) * (1-sin(x)) using the identity: a^2 - b^2 = (a+b) * (a-b).

OpenStudy (anonymous):

ohhhhhh ok :)

OpenStudy (anonymous):

thank you so much! :)

OpenStudy (anonymous):

@alessandro96